Download
recent gain changes n.
Skip this Video
Loading SlideShow in 5 Seconds..
Recent Gain Changes PowerPoint Presentation
Download Presentation
Recent Gain Changes

Recent Gain Changes

0 Views Download Presentation
Download Presentation

Recent Gain Changes

- - - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript

  1. Recent Gain Changes C.Woody HBD Meeting March 23, 2010

  2. Events of March 16th • At around noon on Tuesday March 16th, we got a call from the Counting House that there were no signals in the online monitoring for modules WS1-WS5 and the LV power on the West South preamps was off and would not turn back on. We called John Haggerty and he suspected that the air flow sensor for the WS preamps had sensed a low airflow and had shut them off. • We asked for an immediate access and John, Frank and I went in to investigate, and indeed the airflow sensor had tripped off the power to the WS preamps. However, air was still flowing normally, and Frank made a very slight adjustment to the airflow sensor and the power came back on. Note: there was also a false indication as to which airflow was off. The GUI indicated the ES airflow, but it was in fact the West. This is presumably a GUI problem. Nevertheless, no real airflow cooling was lost to any modules at any time C.Woody, HBD Group Meeting, 3/23/10

  3. Events of March 16th (cont’d) • After the loss of power to the preamps, John did a power cycle of all the LV crates and did a complete download to all the FEMs. Everything came back normally and we finished the access in about half an hour. Note: the HV was set to Standby for the access and the magnets were never switched off since we did not have to move the carriages. • A new fill was obtained around 14:30 and a new run was started (310311). Beginning with this run, the gains in WS2-WS5 were all anomalously high. However, all other gains were normal. C.Woody, HBD Group Meeting, 3/23/10

  4. C.Woody, HBD Group Meeting, 3/23/10

  5. C.Woody, HBD Group Meeting, 3/23/10

  6. C.Woody, HBD Group Meeting, 3/23/10

  7. C.Woody, HBD Group Meeting, 3/23/10

  8. C.Woody, HBD Group Meeting, 3/23/10

  9. Run 310261 These voltages are actually P/T set 2 ! C.Woody, HBD Group Meeting, 3/23/10

  10. Run 310311 Correct voltages for P/T set 3 C.Woody, HBD Group Meeting, 3/23/10

  11. 49.4 42.5 76.7 45.4 75.7 47.2 93.6 49.2 78.4 35.4 C.Woody, HBD Group Meeting, 3/23/10

  12. Scintillation Peaks C.Woody, HBD Group Meeting, 3/23/10

  13. Pedestal Run, Tuesday, Mar 16, after trip, HV at Operational Normal T.Sakaguchi SavePlotrun310328 C.Woody, HBD Group Meeting, 3/23/10

  14. Physics Run, Wednesday, Mar 17, HV at Operational Higher T.Sakaguchi SavePlotrun310382 C.Woody, HBD Group Meeting, 3/23/10

  15. Pedestal Run, Thursday, Mar 18, HV at Operational Normal T.Sakaguchi SavePlotrun310423 C.Woody, HBD Group Meeting, 3/23/10

  16. Test Pulse Run, Thursday, Mar 18, HV at Operational Normal T.Sakaguchi SavePlotrun310422 C.Woody, HBD Group Meeting, 3/23/10

  17. 62.4 GeV C.Woody, HBD Group Meeting, 3/23/10

  18. 62.4 GeV Voltage lowered to 2K setting C.Woody, HBD Group Meeting, 3/23/10

  19. BBC Charge Sum – 62.4 GeV Cut = 20 S.Rolnick C.Woody, HBD Group Meeting, 3/23/10

  20. Gain vs BBC Cut – East – 62.4 GeV S.Rolnick C.Woody, HBD Group Meeting, 3/23/10

  21. Gain vs BBC Cut – West – 62.4 GeV S.Rolnick C.Woody, HBD Group Meeting, 3/23/10

  22. Online vs Offline Gains – Run 310714 – 62.4 GeV I.Raninovich C.Woody, HBD Group Meeting, 3/23/10

  23. Summary • The gain increased suddenly in WS2-5 with no apparent action that would have caused this. • The increase is more than a factor of two in some modules, which is much larger than a P/T correction of one or even two bins. • The HV values are the same as have been used since Jan 30th and we have never seen any increase such as this, even with very large changes in P/T. • The apparent gain increase seems to be confirmed by an increase in the scintillation peaks in central events at 200 GeV C.Woody, HBD Group Meeting, 3/23/10

  24. Summary (cont’d) • There is no observable change in either the pedestal or the test pulse after the trip on March 16th. This would tend to rule out any change in the electronic gain. • The anomalously high gains seem to have been limited to modules WS2-5 during the 200 GeV run. However, other modules in both the east and west are now showing higher gains in the Online Monitoring at 62 GeV. This could be partly due to the BBC cut, even for peripheral events, which show a steep slope of gain vs BBC charge. We need to define a new BBC cut for our Online Monitoring. Offline gains are lower, but do they agree with 200 GeV ? However, WS2-5 still seem to have higher gains than the rest. C.Woody, HBD Group Meeting, 3/23/10